OpenStudy (anonymous):

How does 48 squared equal 4*3^2?

OpenStudy (anonymous):

\[\sqrt{48}= 4\sqrt{3} \]

OpenStudy (anonymous):

\[48=16\times 3\] not \[4\times 9\]

OpenStudy (anonymous):

\[\sqrt{16\times 3}=\sqrt{16}\sqrt3=4\sqrt3\]

OpenStudy (anonymous):

Oh, and so the three stays squared because you can't reduce it or "resquare" it anymore?

OpenStudy (anonymous):

the 3 stays in because 3 is not a perfect square, i.e. the square root of 3 is an irrational number, not an integer

OpenStudy (anonymous):

whereas even i know that the square root of 16 is 4

OpenStudy (anonymous):

Haha, alright, thanks :) One questio though, could you set it up 48 equals 24 x 2?

OpenStudy (anonymous):

Or since both are irrational numbers, you can't set it up that way?

OpenStudy (triciaal):

look at first 2 lines 48 squared is not the same as square root of 48